OpenStudy (anonymous):

f the polynomial x5 − 105 can be split as the product of the polynomials x − 10 and a, what is a?

OpenStudy (anonymous):

\[a \times (x -10) = x^5 - 105\]\[\Rightarrow a = \dfrac{x^5 -105}{x -10}\]do sum ting liek long division.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Just divide \(\Large x^5 - 105\) by \(\Large x - 10\) to get \(\Large a\) \(\Large a=\frac{x^5 - 105}{ x - 10}\) if you meant \(\Large x^5 − 105\), then you can't factor it more.